Hydrofluoric acid dissolves glass by reacting with silica as shown in this balanced chemical equation.

SiO2 + 4 HF → SiF4 + 2 H2O

  1. If 15.75 g SiO2 (molar mass = 60.08 g/mol) and 18.25 g HF (molar mass = 20.01 g/mol) are combined, how many grams of SiF4 (molar mass = 104.1 g/mol) could be produced?
  1. How many grams of SiO2 AND how many grams of HF will remain when the reaction is complete?
  1. If 16.39 g SiF4 were obtained, what was the percent yield?
Answer #1

PART (a):

Given reaction :

SiO2 + 4 HF ------> SiF4 + 2 H2O

No.of moles of SiO2= (15.75 g) / (60.08 g/mol) = 0.262 mole

No.of moles of HF = (18.25) / (20.01 g/mol) = 0.912 mole

From the balanced reaction, we have:

1 mole of SiO2 required 4 moles of HF

Hence, 0.262 mole of SiO2 will require 4*0.262 moles of HF = 1.0486 mole

But, we don't have that much HF, hence, SiO2 will not be completely reacted.

HF will be completely consumed during the reaction. So the limiting reagent is HF.

So the calculations must be based on moles of HF.

From the balanced reaction :

4 mole of HF produces 1 mole of SiF4.

Hence, 1 mole of HF produces 1/4 mole of SiF4.

Therefore, 0.912 mole of HF produces 0.912 *(1/4) mole of SiF4. = 0.228 mole of SiF4.

Therefore, Mass of SiF4 formed = (0.228 mole) x (104.1 g/mol) = 23.74 g

  • Mass of SiF4 formed = 23.74 g

-----------------------------------------------------------------------------

PART (b):

The limiting reagent is HF, Hence HF will be completely consumed during the reaction.

  • Mass of HF remaining = 0 g
  • Mass of SiO2 remaining = (0.262 - 0.912/4) x (60.08 g/mol) = 2.043 g

-----------------------------------------------------------------------------

PART (c):

Percent yield = (actual yield) / (theoretical yield) *100%

Percent yield = (16.39 g) / (23.74 g) *100%

Percent yield = 69.0 %